A 4kg block of ice at -6°C is put in the water (initial temperature 20°C) in an insulated container.

How much water was there initially, if the final temperature of this system after all of the ice melted is 14°C?

Answer in kilograms to the nearest 0.1kg.

following specific heats and latent heats might be useful:

Cwater = 4186J/kgC

Cice = 2090J/kgC

Csteam =1996J/kgC

Lfusion  = 333000J/kg

Lvaporization=2.25x106J/kg


1
Expert's answer
2021-10-07T08:48:23-0400

Given:

"m=4\:\rm kg"

"t_1=-6\:^{\circ}\rm C"

"t_2=20\:^{\circ}\rm C"

"t_3=0\:^{\circ}\rm C"

"t_4=14\:^{\circ}\rm C"

"c_{\rm ice} = 2090\:\rm J/kg
\cdot C"

"c_{\rm water} = 4186\:\rm J/kg
\cdot C"

"\lambda = 333000\:\rm J/kg"

"M-?"


The amount of heat obtained by ice:


"Q_{\rm ice}=Q_1+Q_2+Q_3\\
=c_{\rm ice} m(t_3-t_1)+\lambda m+c_{\rm water}m(t_4-t_3)"

"Q_{\rm ice}=2090*4*6+333000*4+4186*4*14\\
=1616576\:\rm J"

The heat balance equation says

"Q_{\rm ice}=Q_{\rm water}"

Hence


"1616576\:{\rm J}=c_{\rm water}M(t_2-t_4)"

Finally

"M=\frac{1616576}{4186*6}=64.36\:\rm kg"